OpenStudy (anonymous):

Integral x^3 sin x dx

OpenStudy (anonymous):

\[\int\limits_{}^{} x^3 sin x dx\]

OpenStudy (john_es):

Do it by parts. \[\int udv=uv-\int vdu\]

OpenStudy (john_es):

where \[u=x^3\\dv=\sin x dx\\ v=-\cos x\\ du=3x^2 \]You have to redo the process three times.

OpenStudy (anonymous):

\[u=x^3\]\[du = 3x^2 dx\]\[v = -cosx\]\[dv = sinx dx\]

OpenStudy (john_es):

Perfect.

OpenStudy (anonymous):

\[\int\limits_{}^{} x^3 sinx dx = -x^3cosx + \int\limits_{}^{}cosx 3x^2dx\]

OpenStudy (john_es):

Yes, as you see you must repeat the integratin by parts in \[3\int x^2\cos x dx\]

OpenStudy (john_es):

I have to go, but you can check your results here, http://www.wolframalpha.com/input/?i=Integrate%5Bx%5E3+Sin%5Bx%5D%5D

OpenStudy (john_es):

I hope it helps.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[u = x^2\]\[du = 2x dx\]\[v = sinx\]\[dv = cosx dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{} x^2cosxdx = x^2sinx -2\int\limits_{}^{}xsinxdx\]

OpenStudy (anonymous):

Now what?